Fermi National Laboratory
 

Operation Summary for January 7, 2011 - January 10, 2011

 

The Integrated Luminosity for the period from 1/3/11 to 1/10/11 was 10.79 Pb-1 (inverse picobarns). NuMI reported receiving 7.64E18 protons on target during this same period.

Friday Jan 7

The day shift began with the Tevatron (TEV) off for a D-Zero (D0) access, with Operations monitoring the Recycler stash (387.7E10), with the Antiproton Source (Pbar) stacking (83.6E10), with Linac, Booster, and Main Injector (MI) operating normally, with the NuMI, MiniBooNE, and Meson FTBFT-978 (CALICE) experiments taking beam, and with the Collider Detector at Fermilab (CDF) waiting for the next store.

During the day shift, Operations reported that D0 remained in access for the entire shift. CDF accessed for a couple of hours. Cryo system personnel performed maintenance on several wet engines around the TEV. Electrical Support personnel discovered a TEV power supply (HOPS) that had excessive ripple on its output and a blown SCR fuse, which they replaced. Linac personnel replaced the H- Source and began its vacuum pump down.

During the evening shift, D0 came out of access at 6:15 PM. Operators began the TEV checkout. Operations put the TEV into HEP shot setup at 9:06 PM. At 10 PM, the TEV quenched at TEV sectors A1 and D3 just as the ramp reached 980 GeV. Operators consulted with TEV personnel, but they were unable to identify the cause of the quench. Operators began the TEV quench recovery and turn on.

Saturday Jan 8

The midnight shift began with the TEV turn on progressing, with Operations monitoring the stash (306.6E10), with Pbar stacking (16.3E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations put the TEV into shot setup at 2:53 AM. Unfortunately, the TEV quenched at 3:45 AM due to the A2 service building door into the enclosure opening slightly, which dropped the safety system interlocks for A sector. Interlock technicians came in and adjusted the switch on the door. Operators searched and secured TEV sector A and began the TEV quench recovery. Operations established beam to T-978 at 4:45 AM.

During the day shift, the TEV quench recovery ran all shift due to the extensive number of houses that quenched. A TEV quadrupole low beta supply (B0Q6) developed a regulation problem. EE Support personnel came in to investigate.

During the evening shift, Operators prepared D0 for a short access. Operations disabled beam to T-978 at 6 PM. EE Support personnel reported that they had repaired the low beta power supply. At 10:54 PM, the TEV quenched with no beam in the machine. Operators consulted with TEV personnel, but they could not find a cause for the quench. Operators began the TEV recovery.

Sunday Jan 9

The midnight shift began with the TEV recovering from another quench, with Operations monitoring the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations put the TEV into shot setup at 4:24 AM and established store 8410 at 5:23 AM with an average initial luminosity (AIL) of 277.87E30. Unfortunately, the TEV quenched at 5:50 AM due to a B11 vertical separator spark. Operators put the TEV into quench recovery. Operations established beam to T-978 at 6 AM.

During the day shift, Operations kept the TEV in an extended dry squeeze to determine if there were any additional problems with the separator. Operations put the TEV into shot setup at 11:44 AM and established store 8412 at 12:47 PM with an AIL of 307.9E30.

During the evening shift, Operations reported that a Booster power supply tripped off at 4:36 PM due to a relay fault. Booster personnel came in and repaired the supply. The Proton Source resumed sending beam at 6:33 PM. Operations disabled beam to T-978 at 4:36 PM.

Monday Jan 10

The midnight shift began with Operations monitoring store 8412 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 8412 at 3:43 AM, put the TEV into shot setup, and established store 8413 by 4:54 AM with an AIL of 322.28E30. Operators reported replacing a Pbar CAMAC crate power supply that had failed. In the process, 38 mA of antiprotons were lost from the stack. Operations established beam to T-978 at 5:14 AM.

The Plans for Monday

The plans for today are to stack, stash, and store. The Run Coordinator will allow MI personnel to conduct a gap clearing kicker study and allow Recycler personnel to conduct a lifetime study.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 01/10/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory